The Mr.Abdullah has a monthly salary is SR-12,500 & earns from a part time job SR-2,500 / month. He bought a 24 Carat Gold bar of 120gms as an investment an year ago. He also kept 2,500 USD from his annual savings. He also gave his parents SR-2,500pm since the beginning of the year. His house loan is SR-3,000 a month and he purchased 1,500 shares of Al-Rajhi bank as an investment from which the bank gave him SR-5,000 as dividend. Abdullah also won an award and received SR-30,000 Cash prize. His savings are SR-25,000, at the lowest point of year.
Required : How much Zakat should he pay ?
Answer #1

Assuming Value of Gold Nisab per gram is 212.043 SR (current)

Value of Gold Nisab = 120*212.043 = 25445.16

Zakat would be:-

Bank Saving (Lowest point of year) = 25000

Add: Gold value (in Monetary Value) = 25445.16

Add: Money Owed by Mr. Abdullah (Deposits and Loans Made) = 36000

Add: Resale value of shares = Assuming Nil

Total = 86445.16

Zakat = 86445.16*2.5% = SR 2161.13

Thanks and Regards